A box of mass 2.5 kg is pulled up a ramp by a rope at a constant speed. The angle of inclination of the ramp is 15° and the coefficient of kinetic friction of the ramp is 0.28. The rope is parallel to the ramp. Find the tension in the rope. 


1
Expert's answer
2021-10-20T10:41:26-0400

Let's apply the Newton's Second Law of Motion in projections on axis "x"- and "y":


"F_T-mgsin\theta-F_{fr}=0,""N-mgcos\theta=0,""F_T-mgsin\theta-\mu_kN=0,""F_T-mgsin\theta-\mu_kmgcos\theta=0."

From the last equation we can find the tension in the rope:


"F_T=mg(sin\theta+\mu_kcos\theta),""F_T=2.5\ kg\times9.8\ \dfrac{m}{s^2}\times(sin15^{\circ}+0.28\times cos15^{\circ})=13\ N."